A fair coin is tossed. If a head turns up, a fair die is
tossed ; if a tail turns up, 2 fair dice are tossed. What
is the probability that the face (or sum of the faces)
or the die (or dice) is equal to 6?

Answer by Edwin McCravy(20055)

P(T&6 or H&1&5 or H&2&4 or H&3&3 or H&4&2 or H&5&1) =

P(T&6) + P(H&1&5) + P(H&2&4) + P(H&3&3) + P(H&4&2) + P(H&5&1)

P(T&6) = (1/2)(1/6) = 1/12 =  6/72
P(H&1&5) = (1/2)(1/6)(1/6) =  1/72
P(H&2&4) = (1/2)(1/6)(1/6) =  1/72
P(H&3&3) = (1/2)(1/6)(1/6) =  1/72
P(H&4&2) = (1/2)(1/6)(1/6) =  1/72
P(H&5&1) = (1/2)(1/6)(1/6) =  1/72 
----------------------------------
                     Total = 11/72 

Edwin